# journal: TOG

### v1

@@ Abstract.
We introduce MeshPad, a generative approach that creates 3D meshes from sketch inputs. Building on recent advances in artistic-designed triangle mesh generation, our approach addresses the need for interactive artistic mesh creation. To this end, we focus on enabling consistent edits by decomposing editing into ‘deletion’ of regions of a mesh, followed by ‘addition’ of new mesh geometry. Both operations are invoked by simple user edits of a sketch image, facilitating an iterative content creation process and enabling the construction of complex 3D meshes. Our approach is based on a triangle sequence-based mesh representation, exploiting a large Transformer model for mesh triangle addition and deletion. In order to perform edits interactively, we introduce a vertex-aligned speculative prediction strategy on top of our additive mesh generator. This speculator predicts multiple output tokens corresponding to a vertex, thus significantly reducing the computational cost of inference and accelerating the editing process, making it possible to execute each editing step in only a few seconds. Comprehensive experiments demonstrate that MeshPad outperforms state-of-the-art sketch-conditioned mesh generation methods, achieving more than $22\%$ mesh quality improvement in Chamfer distance, and being preferred by $90\%$ of participants in perceptual evaluations.
journal: TOG
[…]
Haoxuan Li
Ziya Erkoc
Lei Li
Daniele Sirigatti
Vladislav Rosov
Angela Dai
Matthias Nießner
@@
Figure 1. MeshPad enables interactive mesh creation and editing with sketches. We decompose this complex task into two sketch-conditioned operations: addition and deletion. Top: our method allows for a user to create and modify artistic-designed triangle meshes by simply drawing and editing 2D sketches, achieving intuitive and interactive 3D modeling. Bottom (left): our method generates a variety of complex yet compact meshes. Bottom (right): our interactive user interface allows users to iteratively edit the mesh, with each edit step taking a few seconds.
1. Introduction
@@
Figure 2. Method Overview. We show the decomposition of mesh creation and editing into addition and deletion operations (left and right, respectively). ADD: in mesh addition, a transformer generates new mesh regions corresponding to newly added strokes (in red) in the input sketch. The generated mesh triangles are then merged with the existing mesh. DEL: in mesh deletion, we show a deletion operation applied to the mesh addition output. We erase sketch strokes to remove corresponding mesh regions, and overlay the erased strokes in red alongside untouched regions to provide more context to the deletion network. We then predict which mesh vertices correspond to regions to be deleted, and prune the mesh accordingly. After each addition or deletion operation, we automatically generate an updated sketch corresponding to the current output mesh, to enable further sketch-based editing.
@@
Triangle meshes are one of the most predominant 3D representations used in 3D production applications, from video games to virtual reality and movies. Mesh creation and editing is thus a central element of computer graphics. In contrast to 3D representations such as voxels (Dai et al., 2017, 2020; Ren et al., 2024), points (Vahdat et al., 2022; Zhou et al., 2021; Yang et al., 2019), or neural implicit representations (Hertz et al., 2022; Mildenhall et al., 2021; Kerbl et al., 2023), triangle meshes represent surfaces in a compact, structured fashion, as well as enabling efficient fine-scale detail and naturally integrating into modern rendering and editing pipelines, achieving high fidelity with relatively few primitives.
@@
Recent advances in generative 3D models have shown significant potential in generating 3D meshes directly (Siddiqui et al., 2023; Chen et al., 2024b, e; Tang et al., 2024; Hao et al., 2024); however, these output meshes are not editable, which is a crucial component for artistic design in content creation. More specifically, artistic content creation is an iterative process that encompasses not only the initial generation, but also requires multiple cycles of manipulation and editing in order to refine an output to achieve a precise artistic vision. In order for such editing to fit with content creation pipelines, various edits must be performed interactively, affecting only the intended region of the mesh.
@@
Algorithmically, MeshPad leverages a hybrid approach combining an autoregressive network for addition and a token-classification network for deletion. Unlike existing artistic-designed mesh generation methods, our addition network generates partial shapes only corresponding to new sketch strokes, rather than repeatedly synthesizing the entire mesh. This naturally solves the problem of preserving unedited mesh regions during editing. Additionally, based on the triangle sequence representation (Chen et al., 2024e), we introduce a vertex-aligned speculative decoder to accelerate autoregressive generation and reduce computational time. Instead of predicting one vertex coordinate at a time, the speculator allows us to predict $3$ coordinates, i.e., one vertex, at a time. We find that jointly training the speculative head and aligning it with the vertex tokens reduces generation time while maintaining generation quality. Furthermore, we implemented a user interface that allows users to interactively edit 3D meshes by drawing 2D sketches. Extensive experiments show that our method outperforms state-of-the-art methods in sketch-conditioned mesh generation, producing cleaner meshes that more accurately align with the sketch input and excelling in partial mesh editing tasks. Finally, our speculative prediction acceleration enables interactive mesh generation and editing within seconds while maintaining overall shape quality.
@@ 2. Related Work
3D Mesh Generation. Direct mesh generation offers the compelling advantage of producing outputs that closely resemble artist-crafted 3D shapes. Early approaches proposed various parameterizations of an irregular mesh structure, including altases (Groueix et al., 2018) and graphs (Dai and Nießner, 2019). Recently, Polygen (Nash et al., 2020) and MeshGPT (Siddiqui et al., 2023) have demonstrated the remarkable potential of transformers in autoregressively generating artistic meshes. Follow-ups such as MeshAnything (Chen et al., 2024b), MeshAnything V2 (Chen et al., 2024e), EdgeRunner (Tang et al., 2024), and Meshtron (Hao et al., 2024) each proposed improvements in tokenization, attention mechanisms, or shape coverage, pushing the limit of direct mesh generation. PolyDiff (Alliegro et al., 2023) instead uses a diffusion backbone for polygonal mesh synthesis. Despite solid progress, these methods focus on full-shape generation and cannot easily handle local editing tasks.
Conditional 3D Generation. The history of sketch-driven methods starts with techniques that inflate 2D contours (Igarashi et al., 1999; Nealen et al., 2007) or retrieve 3D models from sketch-based queries (Eitz et al., 2012; Funkhouser et al., 2003). Modern sketch-based 3D generation methods typically rely on either single-shot generation SDF diffusion models (Zheng et al., 2023; Robin Borth, 2024) or direct editing of neural fields (Mikaeili et al., 2023), which usually struggle to output tidy, artistically crafted meshes.
@@
With the advancement in large text-to-image models, DreamFusion (Poole et al., 2023) proposes score distillation sampling with 2D diffusion for text-to-3D generation, followed by Magic3D (Lin et al., 2023), Fantasia3D (Chen et al., 2023b), and Meta 3D AssetGen (Siddiqui et al., 2024). Text conditions lack precision for fine-grained control (Xu et al., 2023) and additional image condition is often required (Qian et al., 2024; Bala et al., 2024). In contrast, our method uses iterative sketch conditioning for fine-grained shape control.
@@
3D Shape Editing. With the same concept of SDS, many 3D shape editing methods rely on neural representations (Khalid et al., 2025; Liu et al., 2025), and use region-specific masks (Chen et al., 2024a; Barda et al., 2024; Chen et al., 2024d). SPAGHETTI (Hertz et al., 2022), SALAD (Koo et al., 2023), and PartGen (Chen et al., 2024c) advance part-level neural shape editing by enabling disentangled control, diffusion-based manipulation, and multi-view diffusion for generating and editing meaningful 3D parts, respectively.
@@
Recently, sketch-based 3D editing has emerged as an intuitive alternative to text- and image-based methods. SKED (Mikaeili et al., 2023) and SketchDream (Liu et al., 2024) leverage sketch-conditioned 2D diffusion models for 3D modeling. Other approaches, including SENS (Binninger et al., 2024), Doodle your 3D (Bandyopadhyay et al., 2024), and Masked LRM (Gao et al., 2024), focus on abstraction-aware part manipulation, robust part isolation, and masked reconstruction for fast local edits, respectively.
@@
However, such works struggle to produce an artistic mesh. Techniques based on polygonal meshes or other explicit representations can rely on an expensive optimization (Aigerman et al., 2022; Gao et al., 2023), deformation priors (Tang et al., 2022), specialized parametric families (Elrefaie et al., 2024) or procedural generators (Zhao et al., 2024). Such approaches do not generalize easily to arbitrary geometry and have difficulties making local edits without conflicting with the global constraints. By contrast, our approach addresses these issues using a sequence-based mesh representation to perform partial editing by removing and then adding triangles.
Speculative Decoding. Finally, to reduce inference time for our autoregressive network, we adopt a concept of “speculative decoding”, analogous to language Transformers that generate multiple tokens in one pass (Leviathan et al., 2023; Chen et al., 2023a; Miao et al., 2023; Wertheimer et al., 2024). We design a vertex-aligned speculator for the mesh sequence to accelerate partial additions, enabling agile and expressive iterative mesh editing.
3. Method
@@
MeshPad enables interactive mesh creation and editing by iteratively performing addition and deletion operations guided by input sketches. An input sketch image $\mathcal{I}$ guides the generation of a 3D mesh $\mathcal{M}$ as a sequence of mesh triangles. $\mathcal{M}$ can be further edited by removing strokes from $\mathcal{I}$ to delete mesh regions, or adding new strokes to introduce new geometry. To achieve interactive rates for editing, we introduce a vertex-aligned speculator into our mesh addition transformer, significantly accelerating its autoregressive generation. An overview of our approach is shown in Fig. 2.
Figure 3. Model Architecture. We use the Open Pre-trained Transformer (OPT) as the Transformer backbone to process the embedded triangle sequence and the sketch embeddings from a pre-trained image foundation model, RADIO. Top (right): the deletion network is a classification network that labels each input mesh vertex for removal. The deletion head takes the $xyz$ coordinate embeddings of a vertex as input and produces a masking label for the vertex. Bottom (right): the addition network is an autoregressive generation model that predicts new mesh tokens with a speculator. During the speculative prediction, we align the speculator to the vertex axis so that it always takes the $x$ coordinate as input and predicts the vertex’s $y$ and $z$ coordinates. Note that E, M, V, and ¡s¿ represent Transformer hidden states, vertex masks, vertex coordinate tokens, and the ¡split¿ token, respectively.
3.1. Mesh Creation and Editing
@@
We decouple sketch-based mesh creation and editing into two subtasks: addition and deletion. Both mesh operations are conditioned on a sketch image $\mathcal{I}$, a bitmap where colored line strokes indicate regions to be added or deleted. We define a mesh $\mathcal{M}$ as a sequence of triangles ${}\mathcal{M}=\{\mathcal{F}\}$, where =$\mathcal{F}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$ and $\boldsymbol{v}\in\mathbb{R}^{3}$ represents 3D vertex locations.
@@
Line strokes in the input sketch image $\mathcal{I}$ are divided into two mutually exclusive sets, $\mathcal{I}_{k}$ and $\mathcal{I}_{r}$, colored black and red in our visualizations, respectively. When performing an addition or deletion, $\mathcal{I}_{r}$ corresponds to the user edit, i.e., the part to be added or deleted, respectively, while $\mathcal{I}_{k}$ represents the untouched sketch regions. We denote the mesh parts corresponding to $\mathcal{I}_{r}$ and $\mathcal{I}_{k}$ as mutually exclusive sets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, respectively, and thus $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}$. The operation–addition or deletion–is determined by whether the $\mathcal{I}_{r}$ is added as new strokes (addition), or erased from existing strokes (deletion). This also enables simple supervision of both deletion and addition networks by removing or adding back random regions of mesh geometry of 3D shapes, without requiring ground-truth sequences of mesh edits (see Sec. 3.4 for more detail).
@@
(1) $\mathcal{M}^{\prime}_{r}=\left\{\mathcal{F}\in\mathcal{M}|\exists\boldsymbol{v}\in\mathcal{F}:\boldsymbol{v}\in\mathcal{V}^{\prime}_{r}\right\};\quad\mathcal{M}^{\prime}_{k}=\mathcal{M}\setminus\mathcal{M}^{\prime}_{r}.$
@@
(2) $l_{\boldsymbol{v}}=\begin{cases}0&\text{if }\boldsymbol{v}\in\mathcal{V}_{r},\\ 1&\text{otherwise}.\end{cases}$
This deletion is trained with a binary cross-entropy loss between the predicted binary probability per vertex and the vertex label. Once deletion is performed, we assign ${\mathcal{M}\leftarrow\mathcal{M}^{\prime}_{k}}$ and ${\mathcal{M}_{k}\leftarrow\mathcal{M}^{\prime}_{k}}$, and automatically generate the updated $\mathcal{I}_{k}$ corresponding to the new output mesh $\mathcal{M}$ (Sec. 3.4), which can be further edited.
@@
3.2. Network Architecture
The detailed model structure is shown in Fig. 3. We use a pre-trained image encoder RADIO 2.5-h (Ranzinger et al., 2024) to encode the sketch input as tokens. We follow MeshAnythingV2 (Chen et al., 2024e) to use the same Open Pre-trained Transformer (OPT) (Zhang et al., 2022) network as our backbone. We also adopt the tokenizer $\mathcal{T}$ from MeshAnythingV2 to transform a mesh into a sequence $S=\mathcal{T}(\mathcal{M})$. The sequence $S$ is defined as an ordered list consisting of control tokens (¡split¿, ¡start¿, ¡end¿) and vertex coordinate tokens $V$. The ¡start¿ and ¡end¿ tokens indicate the bounds of the sequence. The ¡split¿ token splits the sequence into subsequences containing only vertex tokens, which define a series of adjacent triangles (Chen et al., 2024e).
To perform mesh addition, we use the OPT model to autoregressively predict the next sequence token $S_{r}^{\prime(i+1)}$ given the sketch $\mathcal{I}$, input mesh sequence $S_{k}=\mathcal{T}(\mathcal{M}_{k})$, and the previously generated tokens $S_{r}^{\prime(1\ldots i)}$:
(3) $P\left(S_{r}^{\prime(i+1)}|S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right)=\text{OPT}\left(S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right).$
After generating the sequence $S^{\prime}_{r}$, we obtain an output mesh by detokenizing $S^{\prime}_{r}$ and merging it with the input partial mesh:
(4) $\mathcal{M}^{\prime}=\mathcal{M}_{k}\cup\mathcal{T}^{-1}(S^{\prime}_{r}).$

### v2

# MeshPad: Interactive Sketch-Conditioned Artist-Designed Mesh Generation and Editing

@@ Abstract
We introduce MeshPad, a generative approach that creates 3D meshes from sketch inputs. Building on recent advances in artist-designed triangle mesh generation, our approach addresses the need for interactive mesh creation. To this end, we focus on enabling consistent edits by decomposing editing into ‘deletion’ of regions of a mesh, followed by ‘addition’ of new mesh geometry. Both operations are invoked by simple user edits of a sketch image, facilitating an iterative content creation process and enabling the construction of complex 3D meshes. Our approach is based on a triangle sequence-based mesh representation, exploiting a large Transformer model for mesh triangle addition and deletion. In order to perform edits interactively, we introduce a vertex-aligned speculative prediction strategy on top of our additive mesh generator. This speculator predicts multiple output tokens corresponding to a vertex, thus significantly reducing the computational cost of inference and accelerating the editing process, making it possible to execute each editing step in only a few seconds. Comprehensive experiments demonstrate that MeshPad outperforms state-of-the-art sketch-conditioned mesh generation methods, achieving more than $22\%$ mesh quality improvement in Chamfer distance, and being preferred by $90\%$ of participants in perceptual evaluations.
@@
Figure 1: MeshPad enables interactive mesh creation and editing with sketches. We decompose this complex task into two sketch-conditioned operations: addition and deletion. Top: our method allows for a user to create and modify artist-designed triangle meshes by simply drawing and editing 2D sketches, achieving intuitive and interactive 3D modeling. Bottom (left): our method generates a variety of complex yet compact meshes. Bottom (right): our interactive user interface allows users to iteratively edit the mesh, with each edit step taking a few seconds.
footnotetext: Project page: https://derkleineli.github.io/meshpad/
1 Introduction
@@
Figure 2: Method Overview. We show the decomposition of mesh creation and editing into addition and deletion operations (left and right, respectively). ADD: in mesh addition, a transformer generates new mesh regions corresponding to newly added strokes (in red) in the input sketch. The generated mesh triangles are then merged with the existing mesh. DEL: in mesh deletion, we show a deletion operation applied to the mesh addition output. We erase sketch strokes to remove corresponding mesh regions, and overlay the erased strokes in red alongside untouched regions to provide more context to the deletion network. We then predict which mesh vertices correspond to regions to be deleted, and prune the mesh accordingly. After each addition or deletion operation, we automatically generate an updated sketch corresponding to the current output mesh, to enable further sketch-based editing.
@@
Triangle meshes are one of the most predominant 3D representations used in 3D production applications, from video games to virtual reality and movies. Mesh creation and editing is thus a central element of computer graphics. In contrast to 3D representations such as voxels [19, 20, 49], points [56, 64, 59], or neural implicit representations [28, 42, 31], triangle meshes represent surfaces in a compact, structured fashion, as well as enabling efficient fine-scale detail and naturally integrating into modern rendering and editing pipelines, achieving high fidelity with relatively few primitives.
@@
Recent advances in generative 3D models have shown significant potential in generating 3D meshes directly [51, 16, 17, 55, 27]; however, these output meshes are not editable, which is a crucial component for artistic design in content creation. More specifically, artistic content creation is an iterative process that encompasses not only the initial generation, but also requires multiple cycles of manipulation and editing in order to refine an output to achieve a precise artistic vision. In order for such editing to fit with content creation pipelines, various edits must be performed interactively, affecting only the intended region of the mesh.
@@
Algorithmically, MeshPad leverages a hybrid approach combining an autoregressive network for addition and a token-classification network for deletion. Unlike existing artist-designed mesh generation methods, our addition network generates partial shapes only corresponding to new sketch strokes, rather than repeatedly synthesizing the entire mesh. This naturally solves the problem of preserving unedited mesh regions during editing. Additionally, based on the triangle sequence representation [17], we introduce a vertex-aligned speculative decoder to accelerate autoregressive generation and reduce computational time. Instead of predicting one vertex coordinate at a time, the speculator allows us to predict $3$ coordinates, i.e., one vertex, at a time. We find that jointly training the speculative head and aligning it with the vertex tokens reduces generation time while maintaining generation quality. We further demonstrate our method through a user interface (Fig. 1) that allows users to interactively edit 3D meshes by drawing on 2D sketches. Extensive experiments show that our method outperforms state-of-the-art methods in sketch-conditioned mesh generation, producing cleaner meshes that more accurately align with the sketch input and excelling in partial mesh editing tasks. Finally, our speculative prediction acceleration enables interactive mesh generation and editing within seconds while maintaining overall shape quality.
@@ 2 Related Work
3D Mesh Generation. Direct mesh generation offers the compelling advantage of producing outputs that closely resemble artist-crafted 3D shapes. Early approaches proposed various parameterizations of an irregular mesh structure, including altases [26] and graphs [18]. Recently, Polygen [43] and MeshGPT [51] have demonstrated the remarkable potential of transformers in autoregressively generating artist-crafted meshes. Follow-ups such as MeshAnything [16], MeshAnything V2 [17], EdgeRunner [55], and Meshtron [27] each proposed improvements in tokenization, attention mechanisms, or shape coverage, pushing the limit of direct mesh generation. PolyDiff [3] instead uses a diffusion backbone for polygonal mesh synthesis. Despite solid progress, these methods focus on full-shape generation and cannot easily handle local editing tasks.
@@
Conditional 3D Generation. The history of sketch-driven methods starts with techniques that inflate 2D contours [29, 44] or retrieve 3D models from sketch-based queries [21, 23]. Modern sketch-based modeling methods explored representations like SDF [63, 50], NeRF [41], and CAD commands [35]. With the advancement in large text-to-image models, DreamFusion [45] proposes score distillation sampling with 2D diffusion for text-to-3D generation, followed by Magic3D [37], Fantasia3D [13], and Meta 3D AssetGen [52]. Text conditions lack precision for fine-grained control [58] and additional image condition is often required [46, 4]. In contrast, our method uses iterative sketch conditioning for fine-grained shape control.
@@
3D Shape Editing. With the same concept of SDS, many 3D shape editing methods rely on neural representations [32, 39] and use region-specific masks [15, 6, 11]. SPAGHETTI [28], SALAD [33], and PartGen [12] advance part-level neural shape editing with disentangled control, diffusion-based manipulation, and multi-view diffusion for generating and editing meaningful 3D parts, respectively.
@@
Recently, sketch-based 3D editing has emerged as an intuitive alternative to text- and image-based methods. SKED [41] and SketchDream [38] leverage sketch-conditioned 2D diffusion models for 3D modeling. Other approaches, including SENS [7], Doodle your 3D [5], and Masked LRM [25], focus on abstraction-aware part manipulation, robust part isolation, and masked reconstruction for fast local edits, respectively.
@@
However, such works struggle to produce an artist-designed mesh. Techniques based on polygonal meshes or other explicit representations can rely on an expensive optimization [2, 24], deformation priors [54], specialized parametric families [22] or procedural generators [62]. Such approaches do not generalize easily to arbitrary geometry and have difficulties making local edits without conflicting with the global constraints. In contrast, our approach addresses these issues using a sequence-based mesh representation to perform partial editing by removing and adding triangles.
Speculative Decoding. Speculative decoding accelerates autoregressive generation by predicting multiple tokens in parallel [34, 10, 40, 57]. We extend this approach to mesh sequences with a vertex-aligned speculator, enabling faster partial additions for iterative mesh editing.
3 Method
@@
MeshPad enables interactive mesh creation and editing by iteratively performing addition and deletion operations guided by input sketches. An input sketch image $\mathcal{I}$ guides the generation of a 3D mesh $\mathcal{M}$ as a sequence of mesh triangles. $\mathcal{M}$ can be further edited by removing strokes from $\mathcal{I}$ to delete mesh regions, or adding new strokes for new geometry. To achieve interactive rates for editing, we introduce a vertex-aligned speculator into our mesh addition transformer to accelerate its autoregressive generation. An overview of our approach is shown in Fig. 2.
3.1 Mesh Creation and Editing
@@
We decouple sketch-based mesh creation and editing into two subtasks: addition and deletion. Both operations are conditioned on a sketch image $\mathcal{I}$, a bitmap where colored line strokes indicate regions to be added or deleted. We define a mesh $\mathcal{M}$ as a set of triangles ${}\mathcal{M}=\{\mathcal{F}\}$, where triangles =$\mathcal{F}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$ and $\boldsymbol{v}\in\mathbb{R}^{3}$ represents vertices.
@@
Line strokes in the input sketch image $\mathcal{I}$ are divided into two mutually exclusive sets, $\mathcal{I}_{k}$ and $\mathcal{I}_{r}$, colored black and red in our visualizations, respectively. When performing an addition or deletion, $\mathcal{I}_{r}$ corresponds to the user edit, i.e., the part to be added or deleted, respectively, while $\mathcal{I}_{k}$ represents the untouched sketch regions. We denote the mesh parts corresponding to $\mathcal{I}_{r}$ and $\mathcal{I}_{k}$ as mutually exclusive sets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, respectively, and thus $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}$. The operation—addition or deletion—is determined by whether the $\mathcal{I}_{r}$ is added as new strokes (addition), or erased from existing strokes (deletion). This also enables simple supervision of both deletion and addition networks by removing or adding back random regions of mesh geometry of 3D shapes, without requiring ground-truth sequences of mesh edits (see Sec. 3.4 for more detail).
@@
$\mathcal{M}^{\prime}_{r}=\left\{\mathcal{F}\in\mathcal{M}|\exists\boldsymbol{v}\in\mathcal{F}:\boldsymbol{v}\in\mathcal{V}^{\prime}_{r}\right\};\quad\mathcal{M}^{\prime}_{k}=\mathcal{M}\setminus\mathcal{M}^{\prime}_{r},$ (1)
where $\setminus$ is the set subtraction operator.
@@
$l_{\boldsymbol{v}}=\begin{cases}0&\text{if }\boldsymbol{v}\in\mathcal{V}_{r},\\ 1&\text{otherwise}.\end{cases}$ (2)
The network is trained with a binary cross-entropy loss between the predicted binary probability per vertex and the vertex label. Once deletion is performed, we assign ${\mathcal{M}\leftarrow\mathcal{M}^{\prime}_{k}}$ and ${\mathcal{M}_{k}\leftarrow\mathcal{M}^{\prime}_{k}}$, and automatically generate the updated $\mathcal{I}_{k}$ corresponding to the new output mesh $\mathcal{M}$ (Sec. 3.4), which can be further edited.
@@
Figure 3: Model Architecture. We use the Open Pre-trained Transformer (OPT) as the backbone to process the embedded triangle sequence and the sketch embeddings from a pre-trained image foundation model, RADIO (frozen during training). Top (right): the deletion network is a classification network that labels each input mesh vertex for removal. The deletion head takes the $xyz$ coordinate embeddings of a vertex as input and produces a masking label for the vertex. Bottom (right): the addition network is an autoregressive generation model that predicts new mesh tokens with a speculator. During the speculative prediction, we align the speculator to the vertex axis so that it always takes the $x$ coordinate as input and predicts the vertex’s $y$ and $z$ coordinates. E, M, V, and <s> represent Transformer hidden states, vertex masks, vertex coordinate tokens, and the <split> token, respectively.
3.2 Network Architecture
The detailed model structure is shown in Fig. 3. We use a pre-trained image encoder RADIO 2.5-h [48] to encode the sketch input as tokens. We follow MeshAnythingV2 [17] to use the same Open Pre-trained Transformer (OPT) [61] network as our backbone. We also adopt the tokenizer $\mathcal{T}$ from MeshAnythingV2 to transform a mesh into a sequence $S=\mathcal{T}(\mathcal{M})$. The sequence $S$ is defined as an ordered list consisting of control tokens (<split>, <start>, <end>) and vertex coordinate tokens $V$. The <start> and <end> tokens indicate the sequence bounds. The <split> token splits the sequence into subsequences containing only vertex tokens, which define a series of adjacent triangles [17]. Please refer to the supplemental for more sequence structure details.
To perform mesh addition, we use the OPT model to autoregressively predict the probability distribution $P$ of the next sequence token $S_{r}^{\prime(i+1)}$ given the sketch $\mathcal{I}$, input mesh sequence $S_{k}=\mathcal{T}(\mathcal{M}_{k})$, and the previously generated tokens $S_{r}^{\prime(1\ldots i)}$:
$P\left(S_{r}^{\prime(i+1)}|S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right)=\text{OPT}\left(S_{k},\mathcal{I},S_{r}^{\prime(1\ldots i)}\right).$ (3)
We detokenize the generated sequence $S^{\prime}_{r}$ to mesh triangles and then merge them with the input partial mesh:
$\mathcal{M}^{\prime}=\mathcal{M}_{k}\cup\mathcal{T}^{-1}(S^{\prime}_{r}).$ (4)